For Nickel, the number density is 8×10

23

8×1023 atoms/

cm3 and the electronic configuration is

1s2 2s2 2p6 3s2 3p6 3d8 4s2. The value of saturation magnetisation of Nickel in its ferromagnetic states is : (use

μB= 9.21×10−21 Am2)



1
Expert's answer
2021-10-13T10:42:14-0400
"M_s=0.6\mu_B\frac{\rho N_A}{A_{Ni}}\\
M_s=0.6(4\pi\cdot10^{-7})\frac{(8.9\cdot10^{6})(6.02\cdot10^{23})}{58.7}=5.\cdot10^5\frac{A}{m}"